Let G be the set of positive real numbers except 1. Define α∗β = αlnβ. Then:

(a) show that (G,∗) is a group.

(b) Is G abelian? if not, find its center.

(c) Give an automorphism of G.
1
Expert's answer
2021-07-21T10:56:05-0400

Let "G" be the set of positive real numbers except 1. Define "α∗β = α^{\lnβ}".


(a) Let us show that "(G,\u2217)" is a group. Since "α∗β = α^{\lnβ}>0" and "α^{\lnβ}=1" if and only if "\ln\beta=0"if and only if "\beta=1," we conclude that the operation is defined on the set "G."

Since "(α∗β)*\gamma = α^{\lnβ}*\gamma=(α^{\lnβ})^{\ln\gamma}=α^{\ln\gamma\lnβ}=α^{\lnβ^{\ln\gamma}}
=α^{\ln(β*\gamma)}=α*(β*\gamma)," we conclude that the operation "*" is associative.

Since "α∗e = α^{\ln e}=α= e^{\ln α}=e*α" for any "\alpha\in G," we conclude that "e" is the identity of "G." Taking into account that "e^{\frac{1}{\ln \alpha}}*\alpha=(e^{\frac{1}{\ln \alpha}})^{\ln\alpha}=e^{\frac{\ln\alpha}{\ln \alpha}}=e," we conclude that "\alpha^{-1}=e^{\frac{1}{\ln \alpha}}\in G," and hence "(G,\u2217)" is a group.


(b) Since "α∗β = α^{\lnβ}=β^{\ln α}=β ∗α" for any "α,\beta\in G," we conclude that "G" is abelian group.


(c) For the map "f: G\to G, f(x)=x," we have that "f(x*y)=f(x^{\ln y})=x^{\ln y}=x*y=f(x)*f(y)," and hence "f" is an automorphism of "G".